OpenStudy (anonymous):

input 3,6,9,12 output 9,15,21,27 what function rule produces the out puts

OpenStudy (raden):

let it is the function of f(n) = pn + q when f(3) = 9 it means 3p + q =9 when f(6) = 15 it means 6p + q = 15 subtract both equations above, get 6p + q = 15 3p + q =9 ------------ (-) 3p = 6 p = 6/3 = 2 subtitute p=2 into one of equation, get 3p + q =9 3(2) + q = 9 6 + q = 9 q = 9 - 6 q = 3 so, the general formula of f(n) = pn + q = 2n + 3

OpenStudy (anonymous):

hanks

OpenStudy (raden):

welcome ;)

OpenStudy (anonymous):

:)